# RL-GSBridge: 3D Gaussian Splatting Based Real2Sim2Real Method for Robotic Manipulation Learning

### v1

@@ IV EXPERIMENTS
TABLE IV: Our soft binding constraint reconstruction method compared with GaMeS [9] on multiple foreground objects and background rendering metrics. SSIM is Scale Structural Similarity Index. PSNR is Peak Signal-to-Noise Ratio. LPIPS is Learned Perceptual Image Patch Similarity.
Banana Bear Cake Small-cube Bg (raw) Bg (tablecloth) Bg (plate)
@@ IV-A2 Tasks
IV-A3 Setup
For each task, we conduct both the simulation and the real world. During testing, we divide the 30 × 30 cm section into four quadrants. For each task, we test the policy across a fixed number of positions in each quadrant to calculate the success rate.
@@ IV-A4 Baseline
To demonstrate that our method effectively reduces sim2real gap, we conduct a baseline experiment called RL-sim: using the same learning method but training directly on images rendered from mesh models shaded in the PyBullet simulator. We select two representative and easily shaded objects: the Small-cube and Bear. Grasping experiments for RL-sim are conducted on a foam pad.
@@ IV-B1 Grasping
In Table I, we compare the grasping performance of RL strategies trained with the baseline (RL-sim) and RL-GSBridge in both simulation and real environments. For both the Small-cube with structured geometry and simple textures and the Bear with unstructured geometry and complex textures, RL-sim shows a significant success rate drop when transferring to real scenarios (an average decrease of 80%) due to substantial visual discrepancies. In contrast, RL-GSBridge not only avoids a success rate drop but also improves it in the Bear grasping scenario (an increase of 14.28%). Our analysis suggests that the rough geometry of unstructured objects in simulation (e.g., unintended protrusions or gaps) would increase the difficulty of completing tasks.
Table II shows that RL-GSBridge experiences an average drop of 6.6% of the success rate in sim-to-real transfer across various complex test scenarios, including diverse objects and desktop backgrounds. This demonstrates that the integration of the physic dynamics-based GS rendering effectively bridges the perception gap between simulation and real environments, maintaining stable strategy transfer success rates across a range of scenarios. Additionally, we observe a noticeable decrease in success rates for the Bear in the Tablecloth background scenario, both in simulation and real environments. The primary reason for this drop is the choice of a brown tablecloth, which has similar texture features to the brown toy bear, causing interference with the strategy network’s decision-making process.
@@
IV-B2 Pick-and-Place
As shown in Table III, we test RL-GSBridge’s sim-to-real performance in a pick-and-place task where a cake is placed onto a plate. The results indicate a 4.54% increase in success rate in real environments. Our analysis suggests that this improvement is due to the differences in perception between simulation and reality. In the simulation, even minor excess contacts during the cake placement process are considered as task failures. In contrast, in real environments, some contacts that do not affect the overall task can be tolerated, leading to a higher success rate as the task is ultimately completed successfully.
IV-B3 Comparison of Sim&Real Behavior Consistency
In Fig. 4, we compare the behavior of the robotic arm in the simulator and in real scenarios. With the same initialization settings, RL-GSBridge exhibits behavior highly consistent with the simulation tests during the grasping process, including first-person perspective perception and the arm’s process posture, whereas RL-sim without the GS model shows significant differences.
@@ IV-B4 GS Rendering results of different Mesh Binding approach
In Table IV, we compare the performance of GaMeS [9] and our soft binding constraints GS modeling method under the setting scenarios and achieve the SOTA performance. Furthermore, Fig. 5 shows that our method obtains fewer artifacts and more detailed texture rendering. As a supplement, in Fig. 6, our soft binding constraint GS modeling method achieves consistent results in editing a non-rigid toy bear, which is beneficial for learning manipulation strategies for non-rigid objects.
@@ V CONCLUSIONS
We propose RL-GSBridge, a RL framework for sim-to-real applications. As an attempt to apply the recently successful radiance field reconstruction methods to robotics, RL-GSBridge has shown promising sim-to-real success rates in desktop-level tasks from the robotic arm’s first-person perspective. This motivates us to explore future directions, such as applying and improving our method on more diverse platforms and tackling more complex tasks. For instance, RL-GSBridge currently lacks generalization capabilities, and developing the ability for adaptive learning is a promising area of exploration. We hope RL-GSBridge will encourage more attempts to apply radiance field reconstruction methods in robotics.

### v2

@@ IV EXPERIMENTS
TABLE IV: Our Soft Mesh Binding Method Compared with GaMeS [9] on Multiple Foreground Objects and Background Rendering Metrics. SSIM is Scale Structural Similarity Index. PSNR is Peak Signal-to-Noise Ratio. LPIPS is Learned Perceptual Image Patch Similarity. Bg Refers to the Background, While Contents in Parentheses Explain the Detailed Differences.
Banana Bear Cake Small_cube Bg (Foam Pad) Bg (Tablecloth) Bg (Plate)
@@ IV-A2 Tasks
IV-A3 Evaluation Setup
For each task, we conduct both the simulation and the real world experiments. During testing, we divide the 30 × 30 cm section into four quadrants. For each task, we test the policy across a fixed number of positions in each quadrant to calculate the success rate.
@@ IV-A4 Baseline
To demonstrate that our method effectively reduces Sim2Real gap, we conduct a baseline experiment called RL-sim: using the same learning method but training directly on images rendered from mesh models shaded in the PyBullet simulator. We select two representative and easily shaded objects: the Small_cube and Bear. Grasping experiments for RL-sim are conducted on a foam pad.
@@ IV-B1 Grasping
In Table I, we compare the grasping performance of RL strategies trained with the baseline (RL-sim) and RL-GSBridge in both simulation and real environments. For both the Small_cube with simple geometry and textures and the Bear with complex geometry and textures, RL-sim shows a significant success rate drop when transferring to real world (an average decrease of 80%) due to visual discrepancies. In contrast, RL-GSBridge demonstrates a minor variation in success rates, maintaining high performance as in simulation. Notably, in the Bear grasping scenario, the success rate in the real world has increased by 14.28%. This may be attributed to the suboptimal simulation of Bear’s soft material and non-structured shape in the simulation.
Table II shows that RL-GSBridge experiences an average drop of 6.6% of the success rate in sim-to-real transfer across various complex test scenarios, including diverse objects and desktop backgrounds. This demonstrates that the integration of the physic dynamics-based GS rendering effectively bridges the perception gap between simulation and real environments, maintaining stable strategy transfer across a range of scenarios. Additionally, we observe a noticeable decrease in success rates for the Bear in the Tablecloth background scenario, both in simulation and real environments. The primary reason for this drop is the choice of a brown tablecloth, which has similar texture features to the brown toy bear, causing difficulty in visual perception.
@@
IV-B2 Pick-and-Place
As shown in Table III, we test RL-GSBridge’s Sim2Real performance in pick-and-place task where a cake is placed onto a plate. The results indicate a 4.54% increase of success rate in real environments, mainly due to the differences in physical contacts between simulation and reality. In simulation, even minor excess contacts during the placement process are considered as task failures. In contrast, some contacts in real environments that do not affect the task can be tolerated, leading to better performance since the task is ultimately completed successfully.
IV-B3 Comparison of Sim&Real Behavior Consistency
In Fig. 4, we compare the behavior of the robotic arm in the simulator and in real scenarios. With the same environment, RL-GSBridge exhibits behavior highly consistent with simulation tests during manipulation, whereas RL-sim without the GS model shows significant differences. Notably, blue lights in the camera view of Fig. 4 is caused by the gripper indicator light. The image augmentation during policy training would mitigate this hue effect, ensuring the consistency of policy behavior. Meanwhile, RL-GSBridge sitll ensure the consistency of texture details between simulated and real-world images.
@@ IV-B4 GS Rendering results of different Mesh Binding approach
In Table IV, we compare the performance of GaMeS [9] and our soft mesh binding GS model under various scenarios and achieve the SOTA performance. Furthermore, Fig. 5 shows that our method obtains fewer artifacts and more detailed texture rendering. As a supplement, in Fig. 6, our soft binding constraint GS modeling method achieves consistent results in editing a non-rigid toy bear. Unfortunately, due to the limitations in simulating soft objects in PyBullet, we do not demonstrate comprehensive experiments on deformable objects manipulation. However, this can be explored as a future direction.
@@ V CONCLUSIONS
We propose RL-GSBridge, a real-to-sim-to-real framework for robotic reinforcement learning. As an attempt to apply the recently successful radiance field reconstruction methods to construct a realistic robotic simulator, RL-GSBridge has shown promising sim-to-real success rates in desktop-level tasks. This motivates us to explore future directions, such as investigating the simulation of realistic lighting [40], and integrating RL-GSBridge with advanced large-scale policy models [41] and perception learning methods [42, 43, 44].We hope RL-GSBridge will encourage more attempts to apply radiance field reconstruction methods in robotics.